# Democratizing Text-to-Image Masked Generative Models with Compact Text-Aware One-Dimensional Tokens

### v1

@@ 5.2 Ablation Studies
Table 3: Ablation on Text-Aware De-tokenization Design. To ensure fairness, all models are trained using our improved one-stage training recipe on the DataComp dataset and evaluated in a zero-shot setting on the ImageNet validation set. Relative improvements for text-aware variants are highlighted in blue. #: Number of tokens. c: Channels of continuous tokens.
Effect of Text-Aware De-Tokenization. We evaluate the impact of text-aware de-tokenization (TA-TiTok) in Tab. 3. For consistency, all models are trained using our improved one-stage training recipe. Tokenizers are trained on the DataComp dataset [25] and evaluated in a zero-shot setting on the ImageNet validation set [17], where the caption is simply represented as “A photo of class” without any prompt engineering. We compare two architectures—discrete tokens (VQ) and continuous tokens (KL)—and vary the token count between 32, 64, and 128. As shown in the table, continuous tokens consistently outperform discrete tokens, aligning with findings in [22]. Additionally, TA-TiTok consistently outperforms TiTok (the non-text-aware variant) across all configurations. Notably, the performance gain is most pronounced with a smaller number of tokens (e.g., 32) and with discrete tokens. We attribute this to the fact that the latent tokens primarily capture low-level image details, while high-level semantics provided by text embeddings enrich these representations. Consequently, the improvement is more substantial with fewer tokens (where learning semantic detail is more challenging) and with vector-quantized tokens (where quantization introduces information loss).
[…]
TA-TiTok KL 32 ID 1.62 213.6
Embedding 1.53 222.0
@@
Text Guidance Type. In our text-aware de-tokenization design, we can use either the numerical IDs from the CLIP text tokenizer or the embeddings from the CLIP text encoder. We ablate this design choice in Tab. 4, where the latter option yields a marginal improvement.
5.3 Main Results
MJHQ-30K. We report the zero-shot text-to-image generation results on MJHQ-30K [37] in Tab. 5. MaskGen-XL (discrete tokens) achieves a significantly better FID compared to recent autoregressive models like LlamaGen [57] (7.51 vs. 25.59) and Show-o [63] (7.51 vs. 14.99), both of which also use VQ tokenizers. Additionally, it offers a 18.5$\times$ improvement in inference throughput over Show-o, with further gains (30.3$\times$ faster) achieved by the lighter MaskGen-L model, albeit with a slight performance drop. MaskGen also demonstrates remarkable resource efficiency. MaskGen-L completes training in just 20.0 8-A100 days, while MaskGen-XL finishes within 35.0 8-A100 days, showcasing both strong performance and efficiency.
For continuous tokens, MaskGen delivers competitive results against recent diffusion models. With only 32 tokens, the MaskGen-L outperforms PixArt-$\alpha$ [14] (630M) (7.24 vs. 9.85), offering 1.4$\times$ faster inference throughput while using fewer parameters and requiring less than one fifth of the training resources (18.5 vs. 94.1 8-A100 days). MaskGen-XL further improves the FID score to 6.53, achieving much better performance compared to SDXL [45] (6.53 vs. 8.76), a 2.6B-parameter model trained on high-quality private data, despite MaskGen-XL being trained exclusively on open data for approximately 30.5 8-A100 days.
GenEval. Tab. 6 summaries the zero-shot text-to-image generation results on GenEval [27]. Using discrete tokens, MaskGen-L (568M) achieves an overall score of 0.53, significantly outperforming the recent autoregressive model LlamaGen [57] by 0.21 and performing on par with the larger Show-o [63] (1.3B). Moreover, the larger MaskGen-XL achieves the highest overall score on the benchmark, with a score of 0.57. This result notably surpasses SDXL [45], a 2.6B-parameter model (2.36$\times$ larger than MaskGen-XL) trained on proprietary data.
Meanwhile, our MaskGen-XL with continuous tokens also achieves an overall score of 0.55, comparable to recent diffusion models [50], but with much lower training costs and exclusively trained on open data.
@@ 6 Conclusion
We introduced TA-TiTok, a text-aware 1D tokenizer that improves semantic alignment in text-to-image generation, and MaskGen, a versatile masked generative model supporting both discrete and continuous tokens. Through compact 1D tokenization, MaskGen reduces training costs and accelerates sampling, making efficient, high-quality generation more accessible. By leveraging publicly available data and an efficient tokenizer design, MaskGen democratizes access to high-performance masked generative models. Our release of training code and model weights aims to drive further progress in text-to-image generation.
Limitations and Future Work. While MaskGen achieves competitive generation quality and benchmark scores comparable to recent text-to-image models, including those leveraging proprietary training data, we acknowledge several aspects for future exploration.
Second, the current implementation of MaskGen operates at a resolution of $256\times 256$. However, the scalability of its core architectural design—1-dimensional tokenization and masked generation—has been validated in high-resolution implementations like Muse [12].
@@
  • Sec. B: Ablation studies on the place of text guidance in TA-TiTok.
  • Sec. C: Comparisons between MaskGen using discrete and continuous tokens.
  • Sec. D: Ablation studies on the number of tokens and aesthetic score condition in MaskGen.
@@
  • Sec. G: More qualitative examples generated by MaskGen.
@@ A More Implementation Details
Training Hyper-parameters. Tab. 8 provides the complete list of hyper-parameters used for training MaskGen with both discrete and continuous tokenizers.
@@
Figure 4: Prompts Used for Recaptioning. One of the four prompts are used when recpationing each image. {original_caption} is replaced with the original caption paired with the image.
Figure 5: Re-captioning Results. Augmented captions, generated by Molmo [16], offer richer details and improved alignment with image content.
@@
B Ablation on Text Guidance Place in TA-TiTok
In the design of TA-TiTok, we only incorporate the text guidance (i.e., the text tokens from CLIP) into the tokenizer decoder to better capture high-level semantics and align with textual descriptions during both reconstruction and generation. In this study, we investigate whether injecting text guidance into both the encoder and decoder of TA-TiTok can further enhance the quality of the encoded latent tokens. This evaluation is conducted on the ImageNet validation set [17] using reconstruction metrics in a zero-shot setting, where the caption is represented as “A photo of class” without any prompt engineering. As shown in Tab. 9, injecting textual guidance into both the encoder and decoder has negligible impact on reconstruction quality. This finding suggests that incorporating text guidance in the decoder alone is sufficient to provide semantic information to the model.
MJHQ-30K GenEval
[…]
TA-TiTok VQ Mask. MaskGen-L (ours) 568M 256 $\times$ 256 20.0 30.3 7.74 0.53
TA-TiTok VQ Mask. MaskGen-XL (ours) 1.1B 256 $\times$ 256 35.0 18.5 7.51 0.57
TA-TiTok KL Mask. MaskGen-L (ours) 568M + 44M 256 $\times$ 256 18.5 11.1 7.24 0.52
TA-TiTok KL Mask. MaskGen-XL (ours) 1.1B + 69M 256 $\times$ 256 30.5 9.1 6.53 0.55
Table 10: Comparison of MaskGen using Discrete and Continuous Tokens on MJHQ-30K and GenEval. “VQ” denotes discrete tokenizers and “KL” stands for continuous tokenizers. “type” indicates the generative model type, where “Mask.” refer to masked transformer models. T: Generator training cost, measured in 8 A100 days using float16 precision. I: Generator inference throughput, measured in samples per second on a single A100 with batch size 64 using float16 precision. FID: MJHQ-30K FID. Overall: GenEval average score.
C Comparisons Between MaskGen Using Discrete and Continuous Tokens
Training and Inference Cost Comparisons Between VQ and KL Variants. The VQ variant of MaskGen benefits from faster training and significantly faster inference, primarily due to inherent differences in the diffusion process used in the KL variant. While the KL variant excels in generating more diverse and higher-aesthetic images, this advantage comes with increased computational demands. To address this gap in training and inference efficiency, we employ 128 tokens for the VQ variant and 32 tokens for the KL variant, effectively controlling the training cost to remain at a comparable level, as shown in Tab. 10.
D Ablation Studies for MaskGen
Experimental Setup. For efficient ablation experiments, we utilize the discrete version of MaskGen. Performance is evaluated using the FID metric on MJHQ-30K [37] and the overall GenEval [27] score.
@@ D Ablation Studies for MaskGen
Aesthetic Score Conditioning. Tab. 12 illustrates the impact of incorporating aesthetic score conditioning signals. As shown, using aesthetic scores improves the quality of the generation. In Fig. 6, generated images using different aesthetic score conditioning are presented. Images are generated using the same hyper-parameters and prompts, except for the aesthetic score used for conditioning. The results show that the aesthetic score correlates highly with the dramatic lighting and fine-grained detail. For example, in the third row, we can observe that a higher aesthetic score leads to more detailed depictions of trees and stars in the night sky, while using a low aesthetic score leads to simpler representation. This enables more fine-grained control of the generated images based on the user’s needs.
@@
MJHQ-30K GenEval
[…]
VQ MaskGen-L 64 8.66 0.49
7.85 0.50
Table 12: Zero-Shot Text-to-Image Generation Results on MJHQ-30K and GenEval with Ablation of Additional Conditions. Incorporating aesthetic scores as conditioning signals enhances generation quality.
@@
Table 14: Class-conditional ImageNet-1K $256\times 256$ Generation Results Evaluated with ADM [19], using continuous tokens (i.e., KL architecture). #: Number of tokens. c: Channels of continuous tokens. T: Generator training cost, measured in 8 A100 days using float32 precision. I: Generator inference throughput, measured in samples per second on a single A100 with float32 precision.
@@ F KL variant of TiTok on ImageNet
G Qualitative Examples of MaskGen

### v2

@@ 5.2 Optimized Image Tokenization with TA-TiTok
Table 3: Ablation on Text-Aware De-tokenization Design. All models are trained on DataComp with our improved one-stage recipe and evaluated in a zero-shot setting on the ImageNet validation set. Relative improvements for TA-TiTok are highlighted in blue. #: Number of tokens. c: Channels of continuous tokens.
Effect of Text-Aware De-Tokenization. We evaluate the impact of text-aware de-tokenization (TA-TiTok) in Tab. 3. For consistency, all models are trained using our improved one-stage training recipe. Tokenizers are trained on the DataComp dataset [25] and evaluated in a zero-shot setting on the ImageNet validation set [17], where the caption is simply represented as “A photo of class” without any prompt engineering. We compare two architectures—VQ and KL variants—and vary the token count between 32, 64, and 128. As shown in the table, continuous tokens consistently outperform discrete tokens, aligning with findings in [22]. Additionally, TA-TiTok consistently outperforms TiTok (the non-text-aware variant) across all configurations. Notably, the performance gain is most pronounced with a smaller number of tokens (e.g., 32) and with discrete tokens. We attribute this to the text embeddings supplementing high-level semantic information overlooked by latent tokens. Consequently, the improvement is more substantial with fewer tokens and vector-quantized tokens, where latent tokens experience greater information loss.
@@
5.3 Text-to-Image Generation with MaskGen
MJHQ-30K. We report zero-shot text-to-image generation results on MJHQ-30K [38] in Tab. 4. MaskGen-XL (discrete tokens) achieves a significantly better FID than recent autoregressive models like LlamaGen [62] (7.51 vs. 25.59) and Show-o [68] (7.51 vs. 14.99), both using VQ tokenizers. It also offers 18.5$\times$ higher inference throughput than Show-o, with MaskGen-L being 30.3$\times$ faster, though with a slight performance drop. MaskGen demonstrates impressive resource efficiency, with MaskGen-L completing training in 20.0 8-A100 days and MaskGen-XL in 35.0 8-A100 days.
[…]
GenEval. Tab. 5 summaries the zero-shot text-to-image generation results on GenEval [27]. Using discrete tokens, MaskGen-L (568M) achieves an overall score of 0.53, significantly outperforming the recent autoregressive model LlamaGen [62] by 0.21 and performing on par with the larger Show-o [68] (1.3B). Moreover, the larger MaskGen-XL achieves the highest overall score on the benchmark, with a score of 0.57. This result notably surpasses SDXL [46], a 2.6B-parameter model (2.36$\times$ larger than MaskGen-XL) trained on proprietary data. Meanwhile, our MaskGen-XL with continuous tokens also achieves an overall score of 0.55, comparable to recent diffusion models [54], but with much lower training costs and exclusively trained on open data.
MJHQ-30K GenEval
[…]
TiTok KL MaskGen-L 9.82 0.47
TiTok 8.50 0.49
TA-TiTok 8.13 0.51
TA-TiTok 7.24 0.52
Table 6: Ablation of Text-Aware Design and Aesthetic Score Condition. Incorporating aesthetic scores and exploiting TA-TiTok enhance generation quality.
Effect of Text-Aware Design and Aesthetic Score Conditioning. For efficient ablation studies, we use the continuous version of MaskGen-L. As shown in Tab. 6, adopting TA-TiTok significantly improves generation quality (Row 1 vs. Row 3). Additionally, incorporating aesthetic score conditioning further enhances performance (Row 3 vs. Row 4). Moreover, comparing Row 2 with other configurations suggests that while aesthetic score conditioning helps MaskGen better control image fidelity, its impact is less significant than that of the text-aware tokenizer TA-TiTok.
@@ 6 Conclusion
We introduced TA-TiTok, a text-aware 1D tokenizer that enhances semantic alignment, and MaskGen, a versatile masked generative model supporting discrete and continuous tokens. With compact 1D tokenization, MaskGen lowers training costs and accelerates sampling, enabling efficient, high-quality generation. Leveraging public data, it broadens access to high-performance generative models. All code and model weights is released to promote future research.
Supplementary Material
@@
  • Sec. B: Ablation studies on the type and place of text guidance in TA-TiTok.
  • Sec. C: Ablation studies on the number of tokens and aesthetic score condition in MaskGen.
  • Sec. D: Comparisons between MaskGen using discrete and continuous tokens.
@@
  • Sec. G: More qualitative examples for latent code swapping with TA-TiTok.
  • Sec. H: More qualitative examples generated by MaskGen.
  • Sec. I: Discussions on limitations and future work of TA-TiTok and MaskGen.
@@ A More Implementation Details
Training. We adhere strictly to the hyperparameters used to train TiTok across all TA-TiTok variants. Specifically, TA-TiTok is trained with a batch size of 1024 for 1 epoch (650k steps) on the filtered DataComp dataset, using a maximum learning rate of 1e-4 and a cosine learning rate schedule. For MaskGen with discrete tokens, we employ a batch size of 4096, leveraging weight tying [47] to stabilize training, with a cosine learning rate schedule and a maximum learning rate of 4 × 10-4. For MaskGen with continuous tokens, to accommodate the diffusion loss, we use a constant learning rate schedule with a maximum rate of 1 × 10-4 and a batch size of 2048. Masked tokens are sampled by randomly selecting the masking rate from [0, 1] on a cosine schedule, following MaskGIT [11], and text conditioning is randomly dropped with a 0.1 probability to enable classifier-free guidance [33]. Tab. 8 provides the complete list of hyper-parameters used for training MaskGen with both discrete and continuous tokenizers.
@@
Figure 5: Prompts Used for Recaptioning. One of four prompts is used to recaption each image, where {original_caption} is replaced with the original image caption.
Figure 6: Re-captioning Results. Captions augmented by Molmo [16] offer richer details and improved alignment with image content.
@@
B Ablation Studies for TA-TiTok
Text Guidance Type in TA-TiTok. In our text-aware de-tokenization design, we can use either the numerical IDs from the CLIP text tokenizer or the embeddings from the CLIP text encoder. We ablate this design choice in Tab. 9, where the latter option yields a marginal improvement.
Text Guidance Place in TA-TiTok In the design of TA-TiTok, we only incorporate the text guidance (i.e., the text tokens from CLIP) into the tokenizer decoder to better capture high-level semantics and align with textual descriptions during both reconstruction and generation. In this study, we investigate whether injecting text guidance into both the encoder and decoder of TA-TiTok can further enhance the quality of the encoded latent tokens. This evaluation is conducted on the ImageNet validation set [17] using reconstruction metrics in a zero-shot setting, where the caption is represented as “A photo of class” without any prompt engineering. As shown in Tab. 10, injecting textual guidance into both the encoder and decoder has negligible impact on reconstruction quality. This finding suggests that incorporating text guidance in the decoder alone is sufficient to provide semantic information to the model.
C Ablation Studies for MaskGen
Experimental Setup. For efficient ablation studies, we use the discrete version of MaskGen to analyze the impact of token count. Performance is evaluated using the FID metric on MJHQ-30K [38] and the overall GenEval [27] score. Additionally, we provide visual comparisons to illustrate the effect of aesthetic score conditioning during sampling.
@@ C Ablation Studies for MaskGen
Aesthetic Score Conditioning.  Fig. 7 visualizes images generated with different aesthetic scores while keeping other hyperparameters and prompts constant. The results indicate a strong correlation between higher aesthetic scores and enhanced dramatic lighting and fine-grained details. For instance, in the third row, a higher aesthetic score yields richer depictions of trees and stars in the night sky, whereas a lower score results in simpler representations. This enables precise control over image generation based on user preferences.
D Comparisons Between MaskGen Using Discrete and Continuous Tokens
Training and Inference Cost Comparisons Between VQ and KL Variants. The VQ variant of MaskGen benefits from faster training and significantly faster inference, primarily due to inherent differences in the diffusion process used in the KL variant. While the KL variant excels in generating more diverse and higher-aesthetic images, this advantage comes with increased computational demands. To address this gap in training and inference efficiency, we employ 128 tokens for the VQ variant and 32 tokens for the KL variant, effectively controlling the training cost to remain at a comparable level, as shown in Tab. 4 of the main paper.
@@ D Comparisons Between MaskGen Using Discrete and Continuous Tokens
Table 13: Class-conditional ImageNet-1K $\times 256$ 256 Generation Results Evaluated with ADM [19], using continuous tokens (i.e., KL architecture). #: Number of tokens. c: Channels of continuous tokens. T: Generator training cost, measured in 8 A100 days using float32 precision. I: Generator inference throughput, measured in samples per second on a single A100 with float32 precision.
@@ F KL variant of TiTok on ImageNet
Figure 8: Visualization of Latent Token Attention Map and Latent Code Swapping. The results are from VQ variant of TA-TiTok with 32 tokens. Each latent token attends to prominent semantic and swapping the code leads to appearance changes in the corresponding semantic entity that the latent token focuses on.
G Qualitative Examples of Latent Code Swapping
Fig. 8 provides additional visualizations of latent code swapping in VQ variants of TA-TiTok. The results demonstrate that our proposed 1D tokenization encodes images such that each token captures meaningful semantic elements, enabling image manipulation through latent token swapping without even requiring the generator. This semantically rich tokenization explains why 1D tokenization achieves far more efficient encoding than its 2D counterpart while preserving competitive reconstruction fidelity: the encoder dynamically allocates tokens to perceptually informative regions in the image.
H Qualitative Examples of MaskGen
@@
I Limitations and Future Work
While MaskGen achieves competitive generation quality and benchmark scores comparable to recent text-to-image models, including those leveraging proprietary training data, we acknowledge several aspects for future exploration.
Second, the current implementation of MaskGen operates at a resolution of $\times 256$ 256. However, the scalability of its core architectural design—1-dimensional tokenization and masked generation—has been validated in high-resolution implementations like Muse [12].